OpenStudy (anonymous):

Value of the cot 10 degree

OpenStudy (anonymous):

cot=1/tan

OpenStudy (anonymous):

cot 10=1/tan 10

OpenStudy (anonymous):

cot10=tan^-1(10)=5,67128181961771 (using a calculator)

OpenStudy (anonymous):

\[\cot(x) = \frac{ 1 }{ \tan(x) }\] \[\cot(10) = \frac{ 1 }{ \tan(10) }\] \[\cot(10) = 5.6712818196326647\] \[\cot(10) = 5.67\]

OpenStudy (shiraz14):

@celecity : You're correct

OpenStudy (kc_kennylau):

I think it's finding the exact value?

OpenStudy (kc_kennylau):

If it is the case, \[\cos15^\circ\implies\cos5^\circ\implies\cos10^\circ\implies\sin10^\circ\implies\cot10^\circ\]